X represent a certain number, when the number is multiplied by 3 the result is as adding 34 to the number. What is the number?
Answer by Cromlix(4381)

3x = x + 34
3x - x = 34
2x = 34
x = 17.
Proof:
3(17) = 17 + 34
51 = 51.
Hope this helps.
:-)